### v2

# Genesis: Evolving Attack Strategies for LLM Web Agent Red-Teaming

Abstract
As large language model (LLM) agents automate complex web tasks, they boost productivity while simultaneously introducing new security risks. However, relevant studies on web agent attacks remain limited. Existing red-teaming approaches mainly rely on manually crafted attack strategies or static models trained offline. Such methods fail to capture the underlying behavioral patterns of web agents, making it difficult to generalize across diverse environments. In web agent attacks, success requires the continuous discovery and evolution of attack strategies. To this end, we propose Genesis, a novel agentic framework composed of three modules: Attacker, Scorer, and Strategist. The Attacker generates adversarial injections by integrating the genetic algorithm with a hybrid strategy representation. The Scorer evaluates the target web agent’s responses to provide feedback. The Strategist dynamically uncovers effective strategies from interaction logs and compiles them into a continuously growing strategy library, which is then re-deployed to enhance the Attacker’s effectiveness. Extensive experiments across various web tasks show that our framework discovers novel strategies and consistently outperforms existing attack baselines. Our code is available at https://github.com/CjangCjengh/web_agent_attack.
I Introduction
Large language model (LLM) agents have demonstrated remarkable capabilities across a wide range of web automation tasks, from e-commerce [4, 18, 12, 25] and data collection [5, 3, 19, 7] to complex financial operations [9, 21, 16, 8]. These successes showcase the potential of LLM agents in sequential decision-making and their ability to interact with real-world environments. However, most existing applications focus on task completion and efficiency, while the security implications of granting these agents autonomous access to sensitive resources remain critically underexplored. The autonomy that makes these agents powerful also renders them vulnerable to manipulation through subtle environmental changes. Fig. 1 illustrates a typical scenario where an agent’s intended goal is subverted by a malicious prompt injected into a website.
Figure 1: An example of a web agent attack. The user’s goal is to purchase a computer. However, an attacker injects a hidden instruction into the website’s HTML, commanding the agent to buy an electric fan. The compromised agent disregards the user’s original request and proceeds to buy the fan, demonstrating a successful attack.
Web agent red-teaming is the systematic evaluation of web-based LLM agents through adversarial interactions or attacks, aiming to uncover vulnerabilities and assess their robustness in realistic web environments. Recent work on attacking web agents has made substantial progress. Approaches include injecting adversarial prompts hidden within website HTML content [10, 22] and using reinforcement learning (RL) to optimize these prompts based on the agent’s black-box responses [17].
Nevertheless, the core challenge in web agent red-teaming extends beyond manually crafting adversarial prompts or performing static training. Achieving long-term attack success requires developing a deeper understanding of the web agent’s vulnerabilities. For instance, an attacker should not only recognize instructing an agent to ignore prior commands is effective in a specific case, but also extract, evolve and generalize attack strategies for adaptation across diverse contexts. This strategy-evolving approach is essential for effective red-teaming and comprehensive security analysis; however, such a framework remains unexplored in existing research.
To address this gap, we introduce Genesis, an agentic red‑teaming framework that explicitly involves dynamic evolution of web‑agent attack strategies. Genesis frames red‑teaming as a cycle of strategic learning implemented by three collaborative modules. An Attacker retrieves strategies from the strategy library and performs generic algorithm to generate new strategies and the injection attack, a Scorer evaluates the target web agent’s actions to provide feedback, and a Strategist analyzes the resulting interaction logs to summarize successful patterns into high-level strategies. These strategies are compiled into a continuously growing strategy library to improve the Attacker’s effectiveness.
We validate our approach through extensive experiments against state-of-the-art web agents on a diverse set of real-world web tasks. The results demonstrate that Genesis significantly outperforms baseline methods, achieving higher attack success rates. Furthermore, we show that the strategies discovered by Genesis are novel and transferable to other backend LLMs, highlighting the generalizability of our approach.
Figure 2: The framework of Genesis. It operates as a closed-loop system with three core modules: the Attacker, the Scorer, and the Strategist. The Attacker retrieves relevant strategies from the strategy library, performs generic algorithm to generate new strategies, and then generates the injection payload. The Scorer evaluates the target web agent’s actions to provide feedback. The Strategist analyzes interaction logs to summarize successful patterns into high-level strategies, enriching the continuously growing strategy library, which stores strategies as either natural language descriptions or executable code.
Our contributions can be summarized as follows:
  • To the best of our knowledge, we propose the first agentic red-teaming framework that systematically discovers, summarizes, and evolves web agent attack strategies.
  • We combine the genetic algorithm with a hybrid strategy representation, enabling the evolution of attack strategies to improve web‑agent attack success rates.
  • We demonstrate the effectiveness and generalizability of our framework through comprehensive experiments, showing that Genesis achieves superior performance by discovering novel and transferable attack strategies.
II Related Work
II-A Web Agents
Vision-enabled web agents [24, 2, 23, 11, 19] leverage webpage screenshots alongside or instead of HTML content, achieving substantially improved performance on complex web tasks. These agents typically employ multi-stage reasoning pipelines that combine visual understanding with action planning. While these advances have demonstrated impressive task completion capabilities across e-commerce [4, 18, 12, 25], data collection [5, 3, 19, 7], and other domains, the security implications of granting such agents autonomous access to sensitive web resources remain largely unexplored. Our work specifically targets these state-of-the-art web agents to investigate their vulnerabilities to environmental manipulation attacks.
II-B Red-teaming against Web Agents
Research on adversarial attacks against web agents remains limited, primarily focusing on white-box backdoors [20, 13], visual perturbations [14], or manual content injections [10, 22]. While recent automated methods like AdvAgent [17] employ offline fine-tuning to generate prompts, they produce static attackers incapable of adaptation. This fails to mirror real-world red-teaming, which relies on iterative strategy refinement. We address this gap by proposing an autonomous framework that moves beyond static generation to dynamically discover, summarize, and evolve attack strategies in a fully black-box setting.
III Method
III-A Preliminary
III-A1 Web Agent Formulation
Following Xu et al. [17], a web agent is tasked with executing a sequence of actions $\{a_{1},a_{2},\dots,a_{n}\}$ to fulfill a user’s natural language request on a given website. At each step $t$, the agent’s backend LLM determines the next action $a_{t}$ based on the current observation (which includes the webpage’s HTML content $h_{t}$ and its rendered screenshot $i_{t}$), the task, and the history of previous actions $A_{t-1}=\{a_{1},\dots,a_{t-1}\}$. An action $a_{t}$ is typically a triplet $(o_{t},r_{t},e_{t})$, specifying an operation $o_{t}$ (e.g., “TYPE”), an argument $r_{t}$ (e.g., a string to be typed), and a target HTML element $e_{t}$. The agent interacts with the environment sequentially until the task is completed.
III-A2 Adversarial Setting
Following Xu et al. [17], we operate under a realistic black-box setting. The attacker has no access to the web agent’s internal architecture, model weights, or gradients. The attacker’s capability is limited to modifying the HTML content of the webpage, transforming the benign HTML $h$ into an adversarial version $h_{adv}$. The attack objective is to execute a targeted attack, manipulating the agent into performing a malicious action $a_{adv}=(o_{t},r_{adv},e_{t})$, where the operation and target element $e_{t}$ are unchanged, but the argument $r_{t}$ is replaced with a malicious one, $r_{adv}$. For instance, an agent instructed to purchase stock “A” might be manipulated into purchasing stock “B”.
To remain effective and practical, our attacks must adhere to two critical constraints. First, the modification from $h$ to $h_{adv}$ must not produce any visible changes on the rendered webpage, ensuring that a human user would not detect the manipulation. This is achieved by embedding adversarial payloads in non-rendering HTML attributes or using CSS properties like zero opacity. Second, an attack pattern should be easily retargeted to different malicious arguments without requiring a new optimization process. This is typically achieved by using placeholders within the HTML content that can be deterministically replaced.
III-B Overview
Genesis automates red-teaming by creating a closed-loop system that emulates the cycle of expert learning. A core feature of our framework is its hybrid strategy representation, where strategies are stored as either natural language descriptions or executable code. As shown in Fig. 2, the framework operates as a collaborative system with three core modules that form this iterative learning loop:
  • The Attacker receives a task description, retrieves relevant strategies from a continuously growing strategy library, evolves them using the genetic algorithm, and generates a context-aware environmental injection.
  • The Scorer observes the target web agent’s response to the injection, evaluates the outcome, and assigns a nuanced score reflecting the degree of attack success.
  • The Strategist analyzes the complete interaction log, including the injection and the agent’s behavior, summarizes the underlying principle into a new strategy, and archives it in the strategy library for future use.
This cycle allows Genesis to dynamically discover, refine, and evolve its attack capabilities over time.
III-C Attacker
The Attacker is responsible for crafting and deploying a specific environmental injection for a given task. Its process employs a genetic algorithm to evolve and refine attack strategies from the relevant strategies in the strategy library.
Upon receiving a task description, the Attacker first encodes it into a semantic vector using text-embedding-3-small. This embedding serves as a query to retrieve the top-k most relevant past strategies from the strategy library based on cosine similarity. These retrieved strategies are then evolved based on their past performance scores:
  • Mutation: For strategies with a score below 5, the Attacker applies mutation. It prompts an LLM to randomly modify these less effective strategies, introducing novel variations with the goal of discovering entirely new and potentially more successful attack strategies.
  • Crossover: For all strategies with a score above 5, the Attacker performs crossover. It instructs an LLM to analyze these successful strategies and combine their respective strengths and core principles, thereby generating new, powerful strategies.
The newly generated strategies, along with the retrieved strategies and scores, are used to construct a few-shot prompt for an LLM. This context, comprising both natural language descriptions and code examples from the hybrid strategy library, provides the LLM with rich historical context. This enables it to generate a potent and relevant injection payload for the current task. Inspired by the code-based strategies, the LLM can also generate a Python function to programmatically process and refine the initial injection string, allowing for more complex and structured manipulations. The final output of this process is the injection payload. This payload is then embedded into a placeholder in the benign HTML content $h$ to produce the adversarial version $h_{adv}$. Following the injection, a screenshot of the modified page is rendered. Both the adversarial HTML content $h_{adv}$ and the corresponding screenshot are then provided as input to the target web agent.
III-D Scorer
The Scorer provides the crucial feedback signal that drives the learning process. After the target web agent interacts with the modified environment and completes its actions, the Scorer evaluates the outcome. The scoring mechanism is designed to capture both definitive success and partial progress.
First, it checks for perfect execution: if the agent’s final action $a_{t}$ exactly matches the intended malicious action $a_{adv}$, the attack is considered a complete success and receives a maximum score of 10. In all other scenarios, where the attack fails or only partially succeeds, the Scorer passes the web agent’s full response trace, which includes its reasoning and intermediate actions, along with the attack objective to an LLM. The LLM assesses the agent’s behavior and assigns a nuanced score between 1 and 9, rewarding attempts that manipulate the agent closer to the goal. This feedback is then passed to the Strategist.
TABLE I: Comparison against baselines. We report the attack success rate (ASR), measured as pass@10, on the testing tasks. The evaluation is conducted against two web agents, SeeAct and WebExperT, using three different LLMs as backends.
Method Attack against SeeAct [23] Attack against WebExperT [11]
Finance Medical Housing Cooking Average Finance Medical Housing Cooking Average
GPT-4o as the Backend LLM
GCG [26] 1.1 2.5 2.2 3.4 2.3 0.3 2.0 1.1 2.7 1.5
I-GCG [6] 3.5 3.7 4.2 4.6 4.0 2.9 2.2 3.8 5.5 3.6
AgentAttack [15] 14.7 14.1 15.9 15.8 15.1 13.2 12.8 14.5 13.1 13.4
InjecAttack [22] 24.0 25.3 25.4 26.5 25.3 20.5 20.8 21.3 23.0 21.4
EIA [10] 34.2 34.3 35.5 35.3 34.8 27.9 30.2 31.1 29.5 29.7
AdvAgent [17] 42.8 43.5 44.1 44.0 43.6 37.0 35.9 36.8 36.6 36.6
Genesis (w/o Initialization) 44.1 43.8 43.7 46.3 44.5 37.8 38.0 38.4 38.6 38.2
Genesis 54.1 55.3 43.9 58.7 53.0 46.2 47.5 47.0 49.0 47.4
Gemini-2.5-Flash as the Backend LLM
GCG [26] 1.3 1.5 2.1 2.4 1.8 0.2 1.9 0.5 1.7 1.1
I-GCG [6] 2.4 3.3 3.0 4.1 3.2 2.3 1.2 3.1 3.8 2.6
AgentAttack [15] 11.8 11.4 12.8 12.7 12.2 10.1 10.2 11.9 10.5 10.7
InjecAttack [22] 19.2 20.4 20.5 21.1 20.3 16.4 16.6 17.0 18.4 17.1
EIA [10] 27.5 27.6 28.4 28.2 27.9 22.3 24.2 25.0 23.6 23.8
AdvAgent [17] 34.8 34.3 35.4 35.5 35.0 30.2 28.2 29.9 28.8 29.3
Genesis (w/o Initialization) 34.9 35.7 36.2 37.5 36.1 30.4 30.6 30.7 31.1 30.7
Genesis 43.9 44.0 45.2 46.6 44.9 37.1 38.3 37.6 39.3 38.1
GPT-5 as the Backend LLM
GCG [26] 0.7 0.9 1.5 1.8 1.2 0.2 1.2 0.1 0.8 0.6
I-GCG [6] 1.2 2.0 1.7 2.6 1.9 1.1 0.5 1.9 2.4 1.5
AgentAttack [15] 6.4 6.0 6.9 6.8 6.5 5.3 5.2 6.3 5.5 5.6
InjecAttack [22] 10.1 10.9 10.8 11.4 10.8 8.8 8.8 9.1 9.8 9.1
EIA [10] 14.8 14.6 15.2 15.1 14.9 11.9 12.9 13.4 12.5 12.7
AdvAgent [17] 18.1 24.1 18.6 19.3 20.0 16.0 15.0 16.0 15.3 15.6
Genesis (w/o Initialization) 19.1 18.7 19.7 19.8 19.3 15.8 16.8 16.1 16.8 16.4
Genesis 23.0 23.5 23.8 25.7 24.0 19.7 20.5 20.0 21.1 20.3
III-E Strategist
The Strategist is the cognitive core of Genesis, responsible for learning from experience and expanding the system’s strategic knowledge. It analyzes the logs from each attack attempt to summarize generalizable strategies.
Using an LLM, the Strategist processes the full interaction log, which contains several key components: the task description, the injection used, the agent’s behavior, and the final score. From this log, it summarizes the core principle of the attack into a reusable strategy. The LLM chooses to formulate this strategy as either a natural language description, which captures the conceptual essence of the strategy, or an executable code snippet, which defines a precise, programmatic procedure for generating a similar injection.
Finally, the Strategist archives this new strategy, the original task description, its vector embedding, and the corresponding score into the strategy library. This enriches the library, making the newly discovered strategy available for the Attacker in future rounds and ensuring the system’s knowledge continuously evolves.
TABLE II: Ablation study. We evaluate the ASR of different variants against SeeAct. Genesis represents our full framework.
Variant GPT-4o as the Backend LLM Gemini-2.5-Flash as the Backend LLM
Finance Medical Housing Cooking Average Finance Medical Housing Cooking Average
Genesis (w/o Crossover) 42.5 43.1 35.6 43.8 41.3 35.2 34.8 36.1 34.5 35.2
Genesis (w/o Mutation) 50.3 51.0 40.5 54.2 49.0 41.8 42.1 42.5 41.2 41.9
Genesis (w/o Refinement) 49.0 50.5 39.8 53.6 48.2 40.1 39.8 41.3 42.3 40.9
Genesis (w/o Strategist) 28.1 29.5 30.2 31.8 29.9 23.1 22.8 24.5 25.0 23.9
Genesis (w/o Scorer) 35.7 36.5 30.5 39.2 35.1 29.3 29.9 29.9 30.5 29.8
Genesis (w/o Scorer + Trace) 51.2 52.5 44.2 55.1 50.8 41.5 42.1 42.5 44.0 42.5
Genesis (Text Only) 49.5 51.1 40.1 54.3 48.8 40.1 44.3 41.9 42.5 42.2
Genesis (Code Only) 32.8 34.5 24.9 37.2 32.4 24.8 25.5 26.0 27.6 26.0
Genesis 54.1 55.3 43.9 58.7 53.0 43.9 44.0 45.2 46.6 44.9
III-F Initialization and Evaluation
We first run Genesis on the training tasks to build the strategy library. This pre-learned library is then used to initialize the framework for evaluation on the unseen testing tasks. Since the evaluation proceeds sequentially through the testing tasks, our framework can dynamically learn from each task and continuously enrich its strategy library throughout the process.
IV Experiments
IV-A Setup
IV-A1 Dataset and Metrics
Mind2Web [2] is a benchmark designed for assessing generalist web agents. Following AdvAgent [17], we select a subset of 840 tasks that involve actions with potentially severe consequences. These tasks are drawn from four specific domains: Finance, Medical, Housing, and Cooking. We partition this subset into 240 training tasks and 600 testing tasks for evaluation. For each task, we define a specific attack objective following AdvAgent [17].
Our primary evaluation metric is the attack success rate (ASR), specifically measured as $\mathbf{pass@10}$. For each task, an attack is deemed successful if, within 10 attempts, the action generated by the web agent exactly matches our targeted adversarial action triplet $a_{adv}=(o_{t},r_{adv},e_{t})$. A successful match requires the agent to not only use the malicious argument $r_{adv}$ but also to perform the correct operation $o_{t}$ on the intended HTML element $e_{t}$.
IV-A2 Implementation Details
We compare Genesis against five baseline attack methods: GCG [26], I-GCG [6], AgentAttack [15], InjecAttack [22], EIA [10], and AdvAgent [17]. For baseline methods that require a training phase, we utilize the 240 training tasks for training and evaluate them on the 600 testing tasks. For our Genesis framework, we evaluate two distinct settings. In the first setting, we construct the strategy library by running Genesis on the training tasks. The framework is then initialized with this learned strategy library for evaluation on the testing tasks. In the second setting, Genesis begins with an empty strategy library and progressively builds its knowledge by interacting with the testing tasks. The Attacker retrieves the top-10 most relevant strategies from the strategy library to construct its prompt.
We assess the effectiveness of baselines against two web agents: SeeAct [23] and WebExperT [11]. Both agents are powered by multimodal LLMs (MLLMs). For experiments, we use GPT-4o, Gemini-2.5-Flash [1], and GPT-5 as the backend LLMs for both the target web agents and the attack methods.
IV-B Main Results
Table I presents the results of our evaluation. “Genesis” represents our full framework, which is initialized with a strategy library learned from the training tasks. “Genesis (w/o Initialization)” is a variant that begins without a pre-learned strategy library and builds its strategies progressively during the evaluation on the testing tasks.
Genesis consistently outperforms all baselines. This superiority holds across both target web agents and is evident for all three backend LLMs. The significant performance margin underscores the efficacy of our evolutionary, strategy-driven approach, which moves beyond static optimization to build a dynamic and reusable understanding of agent vulnerabilities.
The comparison between our full framework and Genesis (w/o Initialization) provides crucial insights into the value of the strategy library. The full Genesis, which is initialized with strategies learned from the training tasks, consistently achieves the highest attack success rates. This highlights that the ability to learn, summarize and transfer attack knowledge is a key driver of its success. Notably, even the variant without a pre-learned library demonstrates competitive performance, often surpassing baselines like AdvAgent. This suggests that the framework’s ability to discover and refine strategies dynamically is a powerful mechanism in itself.
IV-C Ablation Study
We systematically remove core modules and functionalities in Genesis to evaluate their impact on the overall ASR. The results are presented in Table II. “w/o Refinement” disables the Attacker’s ability to generate Python functions to refine initial injection strings. “w/o Strategist” removes the Strategist module and the associated strategy library. “w/o Scorer” operates without the Scorer, meaning strategies are stored without performance scores. Consequently, there is no crossover or mutation, and retrieved strategies are added directly to the Attacker’s prompt. In “w/o Scorer + Trace”, we also remove the Scorer but compensate by storing the web agent’s full response trace in the strategy library for use in the Attacker’s few-shot prompts. “Text Only” and “Code Only” restrict the Strategist to representing strategies as either natural language descriptions or executable code, respectively.
The performance drop across all variants compared to the full Genesis framework validates our design choices and demonstrates that each component plays an integral role in the framework’s effectiveness. Removing either the crossover or mutation mechanism also degrades performance, confirming that both components of our genetic algorithm are beneficial for evolving attack strategies.
The most significant performance degradation is observed in the variant without Strategist, which removes the core evolutionary loop and the strategy library. This outcome underscores our central hypothesis: the ability to autonomously discover, summarize, and reuse attack strategies is the primary driver of Genesis’s superior performance. Without this learning mechanism, the attacker’s capability is severely diminished. Furthermore, removing the Scorer also leads to a notable decrease in effectiveness, highlighting the importance of a clear feedback signal. An unevaluated library of strategies is far less useful, as the Attacker cannot distinguish between successful and failed tactics from past attempts.
TABLE III: Cross-model transferability of the strategy library. The notation A $\rightarrow$ B signifies that the strategy library is learned using A as the backend LLM, and subsequently evaluated using B. We report the average ASR against SeeAct.
Transfer Setting Average ASR
GPT-4o $\rightarrow$ GPT-4o 53.0
GPT-4o $\rightarrow$ Gemini-2.5-Flash 41.7
GPT-4o $\rightarrow$ GPT-5 21.5
Gemini-2.5-Flash $\rightarrow$ GPT-4o 54.7
Gemini-2.5-Flash $\rightarrow$ Gemini-2.5-Flash 44.9
Gemini-2.5-Flash $\rightarrow$ GPT-5 20.8
GPT-5 $\rightarrow$ GPT-4o 55.3
GPT-5 $\rightarrow$ Gemini-2.5-Flash 38.2
GPT-5 $\rightarrow$ GPT-5 24.0
IV-D Strategy Transferability
To further assess the generalizability of our approach, we conduct a cross-model transferability experiment. In this setting, we build the strategy library on the 240 training tasks described in Section IV-A1 using one LLM, and then evaluate its effectiveness using a different LLM during the testing phase. The results in Table III reveal that the strategies discovered by Genesis are highly transferable. In nearly all cross-model scenarios, the ASR remains high, significantly outperforming the baseline where no pre-learned library is used in Table I. This indicates that the strategies are not merely exploiting model-specific artifacts but are capturing fundamental, generalizable vulnerabilities inherent in the web agent’s decision-making process.
Figure 3: Case studies of successful attacks. The left case shows a text-based strategy, while the right case shows a code-based strategy, where an executable Python function refines a generated initial injection payload. For brevity and visual clarity, lengthy text within the strategies and payloads has been omitted.
IV-E Case Study
Fig. 3 illustrates two strategies autonomously discovered by Genesis. The first case demonstrates a text-based strategy that devises a complex, mixed-language injection disguised as a high-priority system override. The second case shows a code-based strategy where a Python function programmatically refines a deceptive payload to disrupt input processing, misleading the agent into canceling an appointment. These examples highlight Genesis’s ability to evolve diverse attack patterns, ranging from high-level semantic deception to precise procedural obfuscations.
V Conclusion
In this work, we presented Genesis, a web agent red-teaming framework that systematically discovers, summarizes, and evolves reusable attack strategies. Through comprehensive experiments against state-of-the-art web agents across diverse real-world tasks, we demonstrated that Genesis achieves superior attack success rates while discovering novel and transferable strategies. Our work highlights the importance of strategic summarization in understanding and mitigating vulnerabilities of autonomous agents, providing a foundation for developing more robust and secure web agent systems.
Acknowledgment
This work is supported by the National Natural Science Foundation of China (No. 62406267), Guangdong Provincial Project (No. 2024QN11X072), and Guangzhou Municipal Science and Technology Project (No. 2025A04J4070).